# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ 2 Related Literature
Our paper contributes to the growing bodies of literature on data-driven stochastic systems, manifold learning, and dynamic factor modeling. Diffusion maps were introduced by [coifman2006diffusion] as a geometrically faithful dimensionality reduction technique. [coifman2008diffusion] present diffusion maps for feature extraction in high-dimensional stochastic systems, in particular, the first few eigenfunctions of the backward Fokker–Planck diffusion operator are proposed as a coarse-grained low-dimensional representation for the long-term evolution of a system. Our framework leverages the ADM algorithm first introduced in [singer2008nonlinear], and falls within the larger category of models that find linear reduced-order representations of high-dimensional nonlinear stochastic systems.
@@
ADM recovers intrinsic parametrizations of independent and identically distributed (i.i.d.) data that have undergone a nonlinear transformation. The error analysis of ADM is based on a result from [singer_graph_2006] on the convergence of graph Laplacians, which we generalize from the case of uniform data on a compact manifold to data arising from a time series in Euclidean space. The heart of this analysis relies on a concentration inequality (called “the inequality of Chernoff” in [singer_graph_2006, paragraph before Equation (3.14)]) with the second moment of the summands as variance proxy. In [berry_variable_2016, Appendix B.1], it is noted that Bernstein’s Inequality (see, for instance, [boucheron_concentration_2013, Section 2.8]) is an applicable concentration inequality for the case of i.i.d. samples. In recent years, many concentration inequalities have been generalized for data arising from Markov chains (see, for instance [adamczak_tail_2008, douc_consistency_2011, miasojedow_hoeffdings_2014, paulin_concentration_2015, fan_hoeffdings_2021]), and we will draw on this literature in Section 4.
@@
As noted in the introduction, applying diffusion maps directly to time series data may be problematic since the data is not sampled uniformly and one should ideally take into account the temporal relationships in the data. Other works have also directly tackled the issue of explicitly incorporating the time dependencies into the diffusion coordinates or devised diffusion map-based models for multivariate time series. [shnitzer2020diffusion] propose a nonparametric framework for state-estimation of stochastic dynamical systems that are high-dimensional and evolve according to gradient flows with isotropic diffusion. Their framework combines diffusion maps with Kalman filtering and concepts from Koopman operator theory, and relies on the existence of a linear lifting operator. In Section 3.3, we extend and generalize this framework to admit two datasets, one of responses and one of covariates, to perform supervised learning tasks and ensure the estimated hidden states contain explanatory power for the responses. [shnitzer2022manifold] present a data-driven approach for time series analysis that combines diffusion maps with contracting observers from control theory, and their main goal is to construct a low-dimensional representation of high-dimensional signals generated by dynamical systems without requiring strong modeling assumptions. [lian_multivariate_2015] introduce a method to analyze multivariate time series using diffusion maps on a parametric statistical manifold; the framework handles high-dimensional nonstationary signals that are subject to random interference and ambient noise, which could degrade the performance of more conventional methodologies.
@@
Some of the aforementioned works, for instance, [coifman2008diffusion] and [shnitzer2020diffusion], make the assumption that the dynamics of the intrinsic latent state evolve according to a Langevin equation, a key assumption we carry forward into our work. In those works, it is suggested that the dynamics of the diffusion coordinates should approximately follow linear SDEs. Indeed, in Section 4, we establish a new robustness result showing that the approximation error using linear diffusions remains small when averaging over the data, and provided that some spectral assumptions are satisfied. When it comes to financial applications, it is worth noting that the class of Langevin diffusions encompasses a wide variety of stochastic processes ubiquitous in the finance literature (see Table 2), exemplified by the Ornstein-Uhlenbeck process. For the use of Langevin dynamics to model stock prices see, for instance, the early works [bouchaud1998langevin, Canessa2001].
@@
An important financial application where dynamic factor models can yield powerful methodologies is financial stress testing and scenario analysis, which consists in predicting a portfolio’s profit & loss (P&L) when subsets of the risk factors that drive the portfolio’s value experience shocks. In [haugh2020scenario], scenario analysis is embedded within a dynamic factor framework by defining a linear Gaussian state space model. There, derivatives portfolios are considered and the risk factors are low-dimensional, as is usually the case in stress testing. Our state space model, outlined in 3.3, allows one to define arbitrarily high-dimensional scenarios, and unlike the framework in [haugh2020scenario] is purely data-driven and nonparametric, overcoming restrictive assumptions, like a VAR(1) process for the latent factors. Furthermore, our state space model is explicitly designed to admit two sets of observations (covariates and responses) and considers their joint dependence on the latent state as well as their joint dynamics, thus resulting in a supervised learning dynamic factor framework.
3 The Model
We propose a data-driven approach to dynamic factor modeling and conditional sampling for a supervised learning setting where response variables depend on a high-dimensional set of covariates. Rather than assuming a restrictive parametric model for the dynamics of the observations, we leverage the data-driven anisotropic diffusion maps (ADM) algorithm from [singer2008nonlinear] to learn the dynamics of a set of eigenfunctions associated with an intrinsic latent state. Motivated by the Kalman filtering framework of [shnitzer2020diffusion], we develop a sampling procedure to carry out scenario analysis in the original variable spaces, preserving the explanatory power of the observed data. We adopt the notations found in Table 1 throughout the remainder of the paper. We adopt the following conventions for stochastic processes: $X_{t}$ (upper case Latin letter with subscript) is a continuous time process at time $t$, $x(t)$ (lower case Latin letter with bracketed argument) is a discrete time sample of $X_{t}$, and $x_{t}$ (lower case Latin letter with subscript) is a linear state space approximation for $x(t)$. The convention for Greek letters will depend on context.
@@ 3 The Model
$\nabla$ gradient
$\Delta$ Laplacian
$I$ identity matrix
$C^{2}$ space of twice differentiable functions
$\mathbb{L}^{2}(\mu)$ space of square integrable functions with respect to a measure $\mu$
@@
Table 1: Frequently-used notations.
@@ 3.1 Dynamical System Setting
Given integers $m,n>0$, we model the covariates $X=(X_{t})_{t\geq 0}$ and response variables $Y=(Y_{t})_{t\geq 0}$ as continuous-time stochastic processes taking values in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. We model $X$ and $Y$ as nonlinear mappings of a latent stochastic process $\theta=(\theta_{t})_{t\geq 0}$, which takes values in $\mathbb{R}^{d}$ for some unknown dimension $d$. We introduce our key assumptions next.
@@
where $Z$ is a normalizing constant (see, for instance, [bakry_analysis_2014, Section 1.15.7]). We let $p(w)=\frac{1}{Z}\exp(-U(w))$ denote the density of $\mu$. Condition (3.2) ensures that the Fokker–Planck operator or generator associated with $\theta$ given by
@@
has a discrete spectrum, see [bakry_analysis_2014, Corollary 4.10.9].
@@
Therefore, the spectrum of $-\mathscr{L}$, denoted $0=\lambda_{0}\leq\lambda_{1}\leq\dots$, is contained in $[0,\infty)$. The diffusion maps algorithm uncovers the relationship between the latent variables and the observations by relying on the following fact: any ${f\in\mathbb{L}^{2}(\mu)}$ can be decomposed as
@@
If $\varphi_{k}$ is an eigenfunction of $\mathscr{L}$ with eigenvalue $\lambda_{k}\neq 0$, then equation (3.3) along with a Poincaré Inequality implies that $\lambda_{k}\geq 1/C$. In fact, the above Poincaré inequality is equivalent to the existence of a spectral gap for $-\mathscr{L}$, that is, $\lambda_{1}-\lambda_{0}=\lambda_{1}\geq 1/C$. We remark that the term spectral gap is often employed in the diffusion maps literature to mean a gap in the spectrum other than between $\lambda_{0}$ and $\lambda_{1}$ (see, for instance, [coifman2008diffusion, Section 3.2]).
@@
If $\nabla^{2}U\geq\rho I$ for some $\rho>0$, then $\mu$ satisfies a Poincaré inequality with $C=1/\rho$. For instance, if $U(x)=\frac{1}{2}x^{T}Ax$ with $A$ positive definite, then $\mu$ is a non-degenerate Gaussian, the smallest eigenvalue of $A$ gives the spectral gap $\rho$, and $1/\rho$ is in fact the best Poincaré constant. However, this condition is not necessary, for instance, the two-sided exponential (a.k.a. Laplace) distribution $\frac{1}{2}e^{-|x|}$ on $\mathbb{R}$ satisfies a Poincaré inequality with $C=4$.
@@
where $B^{j}$ is a Brownian motion by Lévy’s Theorem (see, for instance, a similar derivation for the case of Bessel processes in [karatzas_brownian_2014, Proposition 3.3.21]). And similarly, writing $Y=(Y^{1},\dots,Y^{n})$ we have $\mathrm{d}Y_{t}^{j}=\mathscr{L}G(\theta_{t})\mathrm{d}t+\sqrt{2}\nabla G_{j}(\theta_{t})\cdot\mathrm{d}W_{t}$, and so on. These SDEs admit the flexibility to capture many of the standard models of mathematical finance, as shown in Table 2. For illustrative purposes, we will work with a low-dimensional example which gives rise to Ornstein–Uhlenbeck (OU) and Cox–Ingersoll–Ross (CIR) processes:
@@
Suppose that $U$ is given by $U(w)=\frac{1}{2}(w_{1}^{2}+w_{2}^{2})$ for $w\in\mathbb{R}^{2}$. In this case, $\theta$ is a 2-dimensional Ornstein–Uhlenbeck process satisfying $\mathrm{d}\theta_{t}=-\theta_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}$, the Fokker–Planck operator satisfies $\mathscr{L}f(w)=\Delta f(w)-w\cdot\nabla f(w)$, and the invariant distribution $\mu$ is $N(0,I)$. The Hermite polynomials $(h_{k})_{k\geq 0}$ are defined by
@@
The first several Hermite polynomials are given by $h_{0}=1$, $h_{1}(x)=x$, and $h_{2}(x)=\frac{1}{\sqrt{2}}(x^{2}-1)$. Note that $\mathscr{L}h_{k}(w_{1})=-kh_{k}(w_{1})$ and $h_{k}^{\prime}(w_{1})=\sqrt{k}h_{k-1}(w_{1})$ for $w_{1}\in\mathbb{R}$. An orthonormal basis for $\mathbb{L}^{2}(\mu)$ is given by the family $(h_{i,j})_{i,j\geq 0}$ where $h_{i,j}(w):=h_{i}(w_{1})h_{j}(w_{2})$ for any $w\in\mathbb{R}^{2}$. Consider the case where $X_{t}=(\theta_{t}^{1})^{2}+(\theta_{t}^{2})^{2}$ so that
@@
linear exponential Black–Scholes–Merton
@@
Table 2: Common stochastic processes from mathematical finance which can be written in the form $X_{t}=F(\theta_{t})$ with $\mathrm{d}\theta_{t}=-\nabla U(\theta_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}$. These may all be checked using Itô’s formula, see for instance, [jeanblanc_mathematical_2009, Sections 2.6 and 6.4] for the necessary computations in the case of CIR and CEV processes. Note also that a linear $U$ does not satisfy the growth condition (3.2); however, these processes can be approximated locally.
@@
In [singer2008nonlinear], diffusion maps were used to uncover the dynamics of $\theta$ when the components $\theta^{1},\dots\theta^{d}$ are independent homogeneous diffusions of the form $\mathrm{d}\theta_{t}^{i}=b^{i}(\theta_{t}^{i})\mathrm{d}t+\sigma^{i}(\theta^{i})\mathrm{d}W_{t}^{i}$ with drifts $b^{1},\dots,b^{d}$ and non-degenerate diffusion coefficients $\sigma^{1},\dots,\sigma^{d}$. In this case, one may apply a Lamperti Transformation (see [karatzas_brownian_2014, Chapter 5, Exercise 2.20]) to map the diffusion coefficients to constants, provided some integrability and boundedness assumptions hold on $b^{i}$ and $\sigma^{i}$. Specifically, for each $i$, let $g^{i}$ be an antiderivative of $1/\sigma^{i}$ and define $\widetilde{\theta}^{i}_{t}=g^{i}(\theta_{t}^{i})$. Then Itô’s lemma gives
@@
To illustrate, let $w,\widetilde{w}\in\mathbb{R}^{d}$ and set $x=F(w)$ and $\widetilde{x}=F(\widetilde{w})$. Denoting the Jacobian matrix for $F$ by $J_{F}$, a Taylor approximation as in [singer2008nonlinear] gives
@@
The first term is a modified Mahalanobis distance. Hence, distances in the latent space can be approximated using observations of $X$ and/or $Y$ along with estimates of the quadratic covariation processes. In the next subsection, we follow this insight to uncover the eigenvalues and eigenfunctions of the Fokker–Planck operator associated with the dynamics of $\theta$.

### v2

@@ 2 Related Literature
The current paper contributes to the growing bodies of literature on data-driven stochastic systems, manifold learning, and dynamic factor modeling. Diffusion maps were introduced by [14] as a geometrically faithful dimensionality reduction technique. [13] present diffusion maps for feature extraction in high-dimensional stochastic systems, in particular, the first few eigenfunctions of the backward Fokker–Planck diffusion operator are proposed as a coarse-grained low-dimensional representation for the long-term evolution of a system. Our framework leverages the ADM algorithm first introduced in [39], and falls within the larger category of models that find linear reduced-order representations of high-dimensional nonlinear stochastic systems.
@@
In its original setting, ADM recovers intrinsic parametrizations of independent and identically distributed (i.i.d.) data that have undergone a nonlinear transformation. The error analysis of ADM is based on a result from [38] on the convergence of graph Laplacians, which we generalize from the case of uniform data on a compact manifold to data arising from a time series in Euclidean space. The heart of this analysis relies on a concentration inequality (called “the inequality of Chernoff” in [38, paragraph before Equation (3.14) ]) with the second moment of the summands as variance proxy. In [4, Appendix B.1 ], it is noted that Bernstein’s Inequality (see, for instance, [7, Section 2.8 ]) is an applicable concentration inequality for the case of i.i.d. samples. In recent years, many concentration inequalities have been generalized for data arising from Markov chains (see, for instance [1, 18, 31, 32, 19]), and we will draw on this literature in Section 4.
@@
As noted in the introduction, applying diffusion maps directly to time series data may be problematic since the data is not sampled uniformly and one should ideally take into account the temporal relationships in the data. Other works have also directly tackled the issue of explicitly incorporating the time dependencies into the diffusion coordinates or devised diffusion map-based models for multivariate time series. [36] propose a nonparametric framework for state-estimation of stochastic dynamical systems that are high-dimensional and evolve according to gradient flows with isotropic diffusion. Their framework combines diffusion maps with Kalman filtering and concepts from Koopman operator theory, and relies on the existence of a linear lifting operator. In Section 5, we extend and generalize this framework to admit two datasets, one of responses and one of covariates, to perform supervised learning tasks and ensure the estimated hidden states contain explanatory power for the responses. [37] present a data-driven approach for time series analysis that combines diffusion maps with contracting observers from control theory, and their main goal is to construct a low-dimensional representation of high-dimensional signals generated by dynamical systems without requiring strong modeling assumptions. [29] introduce a method to analyze multivariate time series using diffusion maps on a parametric statistical manifold; the framework handles high-dimensional nonstationary signals that are subject to random interference and ambient noise, which could degrade the performance of more conventional methodologies.
@@
Some of the aforementioned works, for instance, [13] and [36], make the assumption that the dynamics of the intrinsic latent state evolve according to a Langevin equation, a key assumption we carry forward into our work. In those works, it is suggested that the dynamics of the diffusion coordinates should approximately follow linear SDEs. Indeed, in Section 4, we establish a new robustness result showing that the approximation error using linear diffusions remains small when averaging over the data, and provided that some spectral assumptions are satisfied. When it comes to financial applications, it is worth noting that the class of Langevin diffusions encompasses a wide variety of stochastic processes ubiquitous in the finance literature (see Table 2), exemplified by the Ornstein-Uhlenbeck process. For the use of Langevin dynamics to model stock prices see, for instance, the early works [6, 8].
@@
An important financial application where dynamic factor models can yield powerful methodologies is financial stress testing and scenario analysis, which consists in predicting a portfolio’s profit & loss (P&L) when subsets of the risk factors that drive the portfolio’s value experience shocks. In [22], scenario analysis is embedded within a dynamic factor framework by defining a linear Gaussian state space model. There, derivatives portfolios are considered and the risk factors are low-dimensional, as is usually the case in stress testing. Our state space model, outlined in Section 5, allows one to define arbitrarily high-dimensional scenarios, and unlike the framework in [22] is purely data-driven and nonparametric, overcoming restrictive assumptions, like a VAR(1) process for the latent factors. Furthermore, our state space model is explicitly designed to admit two sets of observations (covariates and responses) and considers their joint dependence on the latent state as well as their joint dynamics, thus resulting in a supervised learning dynamic factor framework.
3 Preliminaries
In this section, we precisely state our mathematical setting and assumptions (Subsection 3.1), and make explicit the data-driven embedding approach using ADM (Subsection 3.2). Throughout the remainder of the paper, we adopt the notations found in Table 1. In addition, we adopt the following conventions for stochastic processes: $X_{t}$ (upper case Latin letter with subscript) denotes a continuous time process at time $t$, $x(t)$ (lower case Latin letter with bracketed argument) is a discrete time sample of $X_{t}$, and $x_{t}$ (lower case Latin letter with subscript) is a linear state space approximation for $x(t)$. The convention for Greek letters will depend on context.
@@ 3 Preliminaries
$\nabla$ Gradient operator
$\Delta$ Laplacian operator
$C^{2}$ Twice continuously differentiable functions
$\mathbb{L}^{2}(\mu)$ Square-integrable functions w.r.t. measure $\mu$
@@
$\mathscr{L}$ Fokker–Planck generator $\Delta-\nabla U\cdot\nabla$
$\theta_{t}$ Latent state process
$X_{t},Y_{t}$ Covariate and response processes
$\mu$ Invariant measure of $\theta$
$\lambda_{k}$ $k$-th eigenvalue of -$\mathscr{L}$
$\varphi_{k}$ $k$-th eigenfunction of $\mathscr{L}$
$\psi_{k}$ $k$-th diffusion coordinate (empirical eigenvector)
$\mathbf{H}^{x},\mathbf{H}^{y}$ Lifting operators
$\ell$ Dimension of diffusion coordinate space
Table 1: Frequently used notations.
@@ 3.1 Dynamical System Setting
Given integers $m,n>0$, we model the covariates $X=(X_{t})_{t\geq 0}$ and response variables $Y=(Y_{t})_{t\geq 0}$ as continuous-time stochastic processes taking values in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. We model $X$ and $Y$ as nonlinear mappings of a latent stochastic process $\theta=(\theta_{t})_{t\geq 0}$, which takes values in $\mathbb{R}^{d}$ for some unknown dimension $d$. We introduce the key assumptions next.
@@
where $Z$ is a normalizing constant (see, for instance, [2, Section 1.15.7 ]). We let $p(w)=\frac{1}{Z}\exp(-U(w))$ denote the density of $\mu$. Condition (3.2) ensures that the backward Fokker–Planck operator or generator associated with $\theta$ given by
@@
has a discrete spectrum, see [2, Corollary 4.10.9 ], which allows for expansions of $\mathbb{L}^{2}(\mu)$ functions using series of eigenfunctions.
@@
Therefore, the spectrum of -$\mathscr{L}$, denoted $0=\lambda_{0}\leq\lambda_{1}\leq\dots$, is contained in $0,\infty)$. The diffusion maps algorithm uncovers the relationship between the latent variables and the observations by relying on the following fact: any $f\in\mathbb{L}^{2}(\mu)$} can be decomposed as
@@
If $\varphi_{k}$ is an eigenfunction of $\mathscr{L}$ with eigenvalue $\lambda_{k}\neq 0$, then equation (3.3) along with a Poincaré Inequality implies that $\lambda_{k}\geq 1/C$. In fact, the above Poincaré inequality is equivalent to the existence of a spectral gap for -$\mathscr{L}$, that is, $\lambda_{1}-\lambda_{0}=\lambda_{1}\geq 1/C$. We remark that the term spectral gap is often employed in the diffusion maps literature to mean a gap in the spectrum other than between $\lambda_{0}$ and $\lambda_{1}$ (see, for instance, [13, Section 3.2 ]).
@@
If $\nabla^{2}U\geq\rho I$ for some $\rho>0$, then $\mu$ satisfies a Poincaré inequality with $C=1/\rho$. For instance, if $U(x)=\frac{1}{2}x^{T}Ax$ with $A$ positive definite, then $\mu$ is a non-degenerate Gaussian, the smallest eigenvalue of $A$ gives the spectral gap $\rho$, and $1/\rho$ is in fact the best Poincaré constant. However, this condition is not necessary, for instance, the two-sided exponential distribution $\frac{1}{2}e^{-|x|}$ on $\mathbb{R}$ satisfies a Poincaré inequality with $C=4$.
@@
where $B^{j}$ is a Brownian motion by Lévy’s Theorem (see, for instance, a similar derivation for the case of Bessel processes in [26, Proposition 3.3.21 ]). And similarly, writing $Y=(Y^{1},\dots,Y^{n})$ we have $\mathrm{d}Y_{t}^{j}=\mathscr{L}G(\theta_{t})\mathrm{d}t+\sqrt{2}\nabla G_{j}(\theta_{t})\cdot\mathrm{d}W_{t}$, and so on. These SDEs admit the flexibility to capture many of the standard models of mathematical finance, as shown in Table 2. For illustrative purposes, we will work with a low-dimensional example which gives rise to Ornstein–Uhlenbeck (OU) and Cox–Ingersoll–Ross (CIR) processes:
@@
Suppose that $U$ is given by $U(w)=\frac{1}{2}(w_{1}^{2}+w_{2}^{2})$ for $w\in\mathbb{R}^{2}$. In this case, $\theta$ is a 2-dimensional Ornstein–Uhlenbeck process satisfying $\mathrm{d}\theta_{t}=-\theta_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}$, the generator satisfies $\mathscr{L}f(w)=\Delta f(w)-w\cdot\nabla f(w)$, and the invariant distribution $\mu$ is $N(0,I)$. The Hermite polynomials $(h_{k})_{k\geq 0}$ are defined by
@@
The first several Hermite polynomials are given by $h_{0}=1$, $h_{1}(x)=x$, and $h_{2}(x)=\frac{1}{\sqrt{2}}(x^{2}-1)$. Note that $\mathscr{L}h_{k}(w_{1})=-kh_{k}(w_{1})$ and $h_{k}^{\prime}(w_{1})=\sqrt{k}h_{k-1}(w_{1})$ for $w_{1}\in\mathbb{R}$. An orthonormal basis for $\mathbb{L}^{2}(\mu)$ is given by the family $(h_{i,j})_{i,j\geq 0}$ where $h_{i,j}(w):=h_{i}(w_{1})h_{j}(w_{2})$ for any $w=(w_{1},w_{2})\in\mathbb{R}^{2}$. Consider the case where $X_{t}=(\theta_{t}^{1})^{2}+(\theta_{t}^{2})^{2}$ so that
@@
linear exponential geometric Brownian motion
@@
Table 2: Common stochastic processes from mathematical finance which can be written in the form $X_{t}=F(\theta_{t})$ with $\mathrm{d}\theta_{t}=-\nabla U(\theta_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}$. These may all be checked using Itô’s formula, see for instance, [24, Sections 2.6 and 6.4 ] for the necessary computations in the case of CIR and CEV processes. Note also that a linear $U$ does not satisfy the growth condition (3.2); however, these processes can be approximated locally.
@@
In [39], diffusion maps were used to uncover the dynamics of $\theta$ for the case where the components $\theta^{1},\dots\theta^{d}$ are independent homogeneous diffusions of the form $\mathrm{d}\theta_{t}^{i}=b^{i}(\theta_{t}^{i})\mathrm{d}t+\sigma^{i}(\theta^{i})\mathrm{d}W_{t}^{i}$ with drifts $b^{1},\dots,b^{d}$ and non-degenerate diffusion coefficients $\sigma^{1},\dots,\sigma^{d}$. In this case, one may apply a Lamperti Transformation (see [26, Chapter 5, Exercise 2.20 ]) to map the diffusion coefficients to constants, provided some integrability and boundedness assumptions hold on $b^{i}$ and $\sigma^{i}$. Specifically, for each $i$, let $g^{i}$ be an antiderivative of $1/\sigma^{i}$ and define $\widetilde{\theta}^{i}_{t}=g^{i}(\theta_{t}^{i})$. Then Itô’s lemma gives
@@
[…]
@@
The first term on the right is a squared modified Mahalanobis distance. Hence, distances in the latent space can be approximated using observations of $X$ and/or $Y$ along with estimates of the quadratic covariation processes. In the next subsection, we follow this insight to uncover the eigenvalues and eigenfunctions of the Fokker–Planck operator associated with the dynamics of $\theta$.